OpenStudy (anonymous):

find the local max/min/saddle points f(x,y) = (1+xy)(x+y)

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

i can't find the critical points

OpenStudy (anonymous):

i get fx = 1 + 2xy + y^2 and fy = 1 + 2xy + x^2 but how do i solve these?

OpenStudy (zarkon):

you want fx=0 and fy=0 thus fx=fy 1 + 2xy + y^2=1 + 2xy + x^2 y^2=x^2 \(y=\pm x\) so if y=x 1 + 2xx + x^2=0 1+3x^2=0 which is not possible if y=-x \[1 - 2xx + (-x)^2=0\] \[1-2x^2+x^2=0\] \[1-x^2=0\] \[x=\pm 1\] so the solution is (x=1 and y=-1) and (x=-1 and y=1)

OpenStudy (zarkon):

(x=1 and y=-1) or (x=-1 and y=1)

OpenStudy (anonymous):

why can you set fx = fy?

OpenStudy (zarkon):

because they are both equal to zero

OpenStudy (zarkon):

fx=0 fy=0 thus fx=0=fy fx=fy

OpenStudy (anonymous):

okay hang on, let me see if i understand this

OpenStudy (anonymous):

why are the critical points (1,-1) and (-1,1) but not (-1,-1) or (1,1)

OpenStudy (zarkon):

if it were (1,1) or (-1,-1) then y=x...and from above we see that that is not possible

OpenStudy (anonymous):

ahhh okay!

OpenStudy (zarkon):

also try plugging (1,1) into 1 + 2xy + y^2 you get \[1+2+1=3\ne 0\]

OpenStudy (zarkon):

lol 1+2+1=4 which is not 0

OpenStudy (anonymous):

okay can you help me on a similar one?

OpenStudy (anonymous):

i can ask it in a new question